OpenStudy (anonymous):

Fanning and Medal Factor completely: 9x2 + 42x + 49 (3x + 7)(3x − 7) (9x − 7)(9x − 7) (3x + 7)(3x + 7) (9x + 7)(9x + 7)

OpenStudy (anonymous):

9x² + 42x + 49 Find the difference of squares (or the square root of first and last term) 9x²--factor-->3x*3x 49--factor-->7*7 (3x + 7)(3x + 7)

OpenStudy (anonymous):

I just want some one to check for me

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Jamierox4ev3r

mathslover (mathslover):

Good work, you are right @HelloGoodmorning

mathslover (mathslover):

You can check yourself whether you are right or not by : (3x+7)(3x+7) = \((3x+7)^2 \\ \text{As : } \space (a+b)^2 = a^2 + 2ab + b^2 \\ (3x+7)^2 = (3x)^2 + 7^2 + 2(3x)(7) = 9x^2 + 49 + 42 x \\ 9x^2 + 42x + 49 \) Hence verified that you are right!

OpenStudy (anonymous):

Lol thank you very much

mathslover (mathslover):

You're welcome ! :)

OpenStudy (anonymous):

:)